OpenStudy (anonymous):

$15 represents 20% of the spending money Judy is saving for a camp this summer. What is the total amount of money she hopes to take to camp

ILovePuppiesLol (ilovepuppieslol):

lets say that she is saving x dollars so 20% of x is 15 \(\large x \times \large \frac{20}{100}=15\) \(\large x \times \frac{1}{5}=15\) \(\large x=15 \times 5\) \(x=75\)